OpenStudy (anonymous):

Betty had $40 in her pocket. She bought 12 identical bags of popcorn. Betty was left with $10 after buying the popcorn. What was the price, in dollars, of each bag of popcorn? 0.55 2.50 4.17 5.20

OpenStudy (mrnood):

if she started with $40, and was left with $10 then how much did she spend in total?

OpenStudy (anonymous):

huh?

OpenStudy (mrnood):

huh what? if you had $40, and went shopping, when you come back you have $10 left - how much have you spent?

OpenStudy (anonymous):

30 dollars, right

OpenStudy (mrnood):

yes so if you spent $30 on 12 things (for example packets of popcorn) then you have to 'divide ' the 30 dollars ove r12 things so how much did each cost?

OpenStudy (anonymous):

Oh wow so it'll be 2.50! Thanks!

OpenStudy (anonymous):

hey Mr. Nood, can you help me again? Kelso is in charge of a bake sale. On each table, t, there will be a platter of 24 cookies and 2 bowls of brownies with b brownies inside. Kelso needs an additional 150 cookies to replace what he sells. This situation can be written as the following expression: 24t + (2b)t + 150 How many total sweets are there if there are 7 tables and 12 brownies in each bowl? I don't want the answer, but help to figure it out.

OpenStudy (anonymous):

Well I mean I do want the answer, but I'd like to figure it out and learn it as well.

OpenStudy (mrnood):

well first you should start a new thread for a new question but for now: the question tells you that the number of tables is 7 (this is the value for t) the question tells you the number of brownies in the bowl is 12 (this is the value for b) so write the equation \[number of sweets = (24\times t) + (2\times b \times t) +150\] but wherever there is a t put 7 wherever there is a b put 12 then you can work out the answer

OpenStudy (anonymous):

The answer would be 486 if I did it correctly, right? And sorry, I'll make sure to make new questions instead of using this one from now on. Just didn't want to lose your help while I got it. Heh.

OpenStudy (mrnood):

That is the right answer - did oyu use the formula?

OpenStudy (anonymous):

Yessir, I just simply put in the numbers and replaced the symbols.

OpenStudy (anonymous):

Thank you for the help, MrNood. I'm going to close this question and put a new one, it's nice having to learn all this again.